OpenStudy (anonymous):

How do I multiply and express my answer in scientific notation? (7 * 10^4) (2 * 10^2) I got 9 x 10^6 but that's not correct.

hartnn (hartnn):

you are multiplying those numbers, so 7*2 =.... ? its not 9

OpenStudy (anonymous):

14 * 100^8?

hartnn (hartnn):

why 100 now ? the 10^6 part was correct.

hartnn (hartnn):

\(10^4.10^2 = 10^{4+2}=10^6 \)

OpenStudy (anonymous):

oh alright

OpenStudy (anonymous):

1.4 x 10^6 is the answer right?

hartnn (hartnn):

you got 14 *\(10^6= 1.4*10^7\) did you get this?

OpenStudy (anonymous):

yeah I get it

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

But what about significant figures @hartnn

hartnn (hartnn):

the what ?

OpenStudy (anonymous):

Ok! it's right.

OpenStudy (anonymous):

Sorry.